# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ A.2 Alternative Constraints for the EF- and EEF-based Models
First, in our version of EF-CYCLE we included constraints (19). These constraints, which are based on the set $\mathcal{P}_{K-1}$ of maximal cycle-feasible paths, forbid all cycles with length exceeding $K$. Alternatively, in the original version of EF-CYCLE that was proposed by Abraham et al., (2007) and Roth et al., (2007), different long-cycle elimination constraints were used. These are based on the set $\mathcal{P}_{K}$ of minimal cycle-infeasible paths, which is defined like $\mathcal{P}_{K-1}$, but where each path has length $K$ instead of $K-1$. Note that a feasible solution can contain at most $K-1$ arcs per minimal cycle-infeasible path, leading to the following constraints:
@@
Moreover, in EF-CHAIN-EXP and EEF-CHAIN-EXP, we considered constraints (26) and (50), which exclude all cycles of length at most $L-1$ and $L-2$, respectively. As an alternative, Anderson et al., (2015) (see their constraints [7] and [S12]) proposed the following “cut set” constraints that can be used in EF-CHAIN-EXP:
@@
Finally, in our version of EEF-CYCLE we included constraints (34), which simultaneously exclude cycles of length exceeding $K$ and reduce symmetry by requiring that if any arcs are selected in some subgraph $G^{s}$ then at least one such arc must leave vertex $s$. Alternatively, in the original version of EEF-CYCLE that was proposed by Constantino et al., (2013), the following two sets of constraints were used instead:
@@
which separately forbid long cycles and break symmetry, respectively. However, our preliminary computational experiments showed that our new combined version of the constraints resulted in more optimal solutions and smaller average running times. Furthermore, note that in theory the LP-relaxation of the model can be improved by including constraints (60) on top of constraints (34), but our experiments showed that the resulting model had a worse computational performance overall than the model with only the latter constraints. Similarly, constraints (49) in EEF-CHAIN-EXP and EEF-CHAIN-MTZ could also be replaced by the following two separate sets of constraints:
@@
as proposed in the original version of the model by Anderson et al., (2015). However, our new version performed better in our preliminary experiments.
A.3 EF-HYBRID
@@
Let $\mathcal{C}_{>K}$ be the set of cycles of length more than $K$. Moreover, let $\mathcal{P}^{*}_{L-1}\subseteq\mathcal{P}^{\prime}_{L-1}$ be the subset of minimal chain-infeasible paths $p$ for which the first vertex $p_{1}$ is adjacent to at least one NDD $v\in\mathcal{N}$. Then, defining binary decision variables $z_{uv}$ for every arc $(u,v)\in\mathcal{A}$, taking value $1$ if arc $(u,v)$ is selected, and value 00 otherwise, we obtain the following model:
[…]
@@
Constraints (64) forbid all cycles of length more than $K$, whereas constraints (65) forbid all chains of length more than $L$. Constraints (64) were also used in the model by Anderson et al., (2015), whereas constraints (65) are an improved version of constraints proposed by Constantino et al., 2013).
Example 1.
(continued) Consider the instance presented in Figure 3 for $K=2$ and $L=3$. Like EF-CHAIN-EXP, EF-HYBRID contains $14$ variables, one for each arc in $\mathcal{A}$. For instance, the solution consisting of cycle $\left\langle 1,4,1\right\rangle$ and chain $\left\langle 6,2,3,\tau\right\rangle$ is obtained by setting $z_{14}=z_{41}=z_{62}=z_{23}=z_{3\tau}=1$. Several long cycles and long chains must be forbidden. For example, cycle $\left\langle 2,3,4,2\right\rangle$ is excluded by the constraint (64) corresponding to this cycle, and chain $\left\langle 5,1,2,3,\tau\right\rangle$ is excluded by constraint (65) for minimal chain-infeasible path $\left\langle 1,2,3\right\rangle$.
Regarding model-related improvements, we can remove all vertices and arcs that cannot appear in any cycle of length at most $K$ or any chain of length at most $L$, which we elaborate on in Appendix A.5. Furthermore, it is also possible to omit the variables $z_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$, as described in Appendix A.6. Moreover, note that the number of constraints in EF-HYBRID is exponential in both $K$ and $L$, due to constraints (64) and (65). However, these constraints can be handled through constraint generation.
@@
A.4 Summary of model properties
@@
Cycles CF-CYCLE $O(|\mathcal{R}|^{K})$ $O(|\mathcal{R}|)$
HCF-CYCLE $O(|\mathcal{R}|^{1+\left\lceil K/2\right\rceil})$ $O(|\mathcal{R}|^{2})$
@@
EEF-CYCLE $O(|\mathcal{R}||\mathcal{A}_{\mathcal{R}}|)$ $O(|\mathcal{R}|^{2})$
PIEF-CYCLE $O(K|\mathcal{R}||\mathcal{A}_{\mathcal{R}}|)$ $O(K|\mathcal{R}|^{2})$
Chains CF-CHAIN $O(|\mathcal{N}||\mathcal{R}|^{L-1})$ $O(|\mathcal{N}|+|\mathcal{R}|)$
HCF-CHAIN $O(|\mathcal{N}||\mathcal{R}|^{\left\lfloor L/2\right\rfloor}+|\mathcal{R}|^{\left\lceil L/2\right\rceil})$ $O(|\mathcal{N}|+|\mathcal{R}|)$
@@
EF-CHAIN-MTZ $O(|\mathcal{A}|)$ $O(|\mathcal{N}|+|\mathcal{A}_{\mathcal{R}}|)$
@@
EEF-CHAIN-MTZ $O(|\mathcal{N}||\mathcal{A}|)$ $O(|\mathcal{N}||\mathcal{R}|+|\mathcal{A}_{\mathcal{R}}|)$
PIEF-CHAIN $O(L|\mathcal{A}|)$ $O(|\mathcal{N}|+L|\mathcal{R}|)$
@@
A.5 Graph Reduction Algorithms
@@
EF-CYCLE ~=$\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:\exists u\in\mathcal{R}\text{ s.t. }(u,v)\in\mathcal{A}_{\mathcal{R}},d_{vu}+1\leq K\}$
@@
EF-CHAIN-EXP/MTZ ~=$\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:d_{\mathcal{N}v}+1\leq L\}$
@@
EF-HYBRID ~=$\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:(\exists u\in\mathcal{R}\text{ s.t. }(u,v)\in\mathcal{A}_{\mathcal{R}},d_{vu}+1\leq K)\vee(d_{\mathcal{N}v}+1\leq L)\}$
@@
$\mathcal{K}^{s}(u,v)=\begin{cases}\{1\}&\text{ if }u=s\text{ and }d^{s}_{vs}\leq K-1,\\ \emptyset&\text{ if }u=s\text{ and }d^{s}_{vs}>K-1,\\ \{k\in\{2,\ldots,K\}:d^{s}_{su}\leq k-1\}&\text{ if }v=s,\\ \{k\in\{2,\ldots,K-1\}:d^{s}_{su}\leq k-1,d^{s}_{vs}\leq K-k\}&\text{ otherwise},\end{cases}$
@@
Instead, we propose to construct the sets $\mathcal{K}^{s}(u,v)$ by running Algorithm 1 once for each subgraph $\mathcal{G}^{s}$ for $s\in\mathcal{R}$. Note that every run is essentially a breadth-first search of the considered subgraph $\mathcal{G}^{s}$, starting from vertex $s$, where a position $k$ is added to the set of possible positions for an arc if (i) the tail of the arc is the head of at least one arc for which position $k-1$ is possible, and (ii) the cycle can still be closed using at most $K-k$ arcs.
@@
1:$\mathcal{K}^{s}(u,v)\leftarrow\emptyset$ for $(u,v)\in\mathcal{A}^{s}$;
2:$\mathcal{S}\leftarrow\{s\}$
@@
1:$\mathcal{K}^{\prime}(u,v)\leftarrow\emptyset$ for $(u,v)\in\mathcal{A}$;
2:$\mathcal{S}\leftarrow\mathcal{N}$
@@
A.6 Implicitly Dealing with the Terminal Vertex

### v2

@@ Appendix B Alternative Constraints for the EF- and EEF-based Models
First, in our version of EF-CYCLE we included constraints (19). These constraints, which are based on the set $\mathcal{P}_{K-1}$ of maximal cycle-feasible paths, forbid all cycles with length exceeding $K$. Alternatively, in the original version of EF-CYCLE that was proposed by Abraham et al., (2007) and Roth et al., (2007), different long-cycle elimination constraints were used. These are based on the set $\mathcal{P}_{K}$ of minimal cycle-infeasible paths, which is defined like $\mathcal{P}_{K-1}$, but where each path has length $K$ instead of K1K-1. Note that a feasible solution can contain at most K1K-1 arcs per minimal cycle-infeasible path, leading to the following constraints:
@@
Moreover, in EF-CHAIN-EXP and EEF-CHAIN-EXP, we considered constraints (26) and (50), which exclude all cycles of length at most L1L-1 and L2L-2, respectively. As an alternative, Anderson et al., 2015b (see their constraints [7] and [S12]) proposed the following “cut set” constraints that can be used in EF-CHAIN-EXP:
@@
Finally, in our version of EEF-CYCLE we included constraints (34), which simultaneously exclude cycles of length exceeding $K$ and reduce symmetry by requiring that if any arcs are selected in some subgraph $\mathcal{G}^{s}$ then at least one such arc must leave vertex $s$. Alternatively, in the original version of EEF-CYCLE that was proposed by Constantino et al., (2013), the following two sets of constraints were used instead:
@@
which separately forbid long cycles and break symmetry, respectively. However, our preliminary computational experiments showed that our new version of the model with the combined constraints resulted in more instances being solved to optimality and smaller average running times. Furthermore, note that in theory the LP-relaxation of the model can be improved by including constraints (60) on top of constraints (34), but our experiments showed that the resulting model had a worse computational performance overall than the model with only the latter constraints. Similarly, constraints (49) in EEF-CHAIN-EXP and EEF-CHAIN-MTZ could also be replaced by the following two separate sets of constraints:
@@
as proposed in the original version of the model by Anderson et al., 2015b . However, our new version performed better in our preliminary experiments.
Appendix C The EF-HYBRID Model
@@
Let $\mathcal{C}_{>K}$ be the set of cycles of length more than $K$. Moreover, let $\mathcal{P}^{*}_{L-1}\subseteq\mathcal{P}^{\prime}_{L-1}$ be the subset of minimal chain-infeasible paths $p$ for which the first vertex $p_{1}$ is adjacent to at least one NDD $v\in\mathcal{N}$. Then, defining binary decision variables $z_{uv}$ for every arc $(u,v)\in\mathcal{A}$, taking value $1$ if arc $(u,v)$ is selected, and value 0 otherwise, we obtain the following model:
$(\texttt{EF-HYBRID})\quad$ $\max$ $\sum_{(u,v)\in\mathcal{A}}w_{uv}z_{uv}$ (63)
@@
Constraints (64) forbid all cycles of length more than $K$, whereas constraints (65) forbid all chains of length more than $L$. Constraints (64) were also used in the model by Anderson et al., 2015b , whereas constraints (65) are an improved version of constraints proposed by Constantino et al., 2013).
Regarding model-related improvements, we can remove all vertices and arcs that cannot appear in any cycle of length at most $K$ or any chain of length at most $L$, which we elaborate on in Appendix F. Furthermore, it is also possible to omit the variables $z_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$, as described in Appendix G. Moreover, note that the number of constraints in EF-HYBRID is exponential in both $K$ and $L$, due to constraints (64) and (65). However, these constraints can be handled through constraint generation.
@@
Appendix D Illustrative example
In Figure 3, we provide an example instance of KE-Opt which we subsequently use to illustrate all considered models.
Figure 3: KE-Opt instance with optimal solution for $K=2$ and $L=3$.
[…]
Introduction to instance:
This instance has $\mathcal{R}=\{1,2,3,4\}$ and $\mathcal{N}=\{5,6\}$, and all arcs have unitary weight (i.e., we consider the unweighted cycles-and-chains case). An optimal solution for $K=2$ and $L=3$ is depicted in bold. This solution consists of cycle $\left\langle 1,4,1\right\rangle$ of length $2$, chain $\left\langle 5,\tau\right\rangle$ of length $1$, and chain $\left\langle 6,2,3,\tau\right\rangle$ of length $3$, resulting in a total of $6$ transplants.
CF-CYCLE:
Suppose that $K\in\{2,3,4\}$. Set $\mathcal{C}_{\leq K}$ only consists of cycle $\left\langle 1,4,1\right\rangle$ if $K=2$, also contains cycle $\left\langle 2,3,4,2\right\rangle$ if $K=3$, and additionally also cycle $\left\langle 1,2,3,4,1\right\rangle$ if $K=4$. That is, when $K=4$, CF-CYCLE contains $3$ variables, one for each cycle of length at most $4$.
CF-CHAIN:
Suppose that $L=3$. We have $\mathcal{C}^{\prime}_{\leq L}=\{\left\langle 5,\tau\right\rangle,\left\langle 6,\tau\right\rangle,\left\langle 5,1,\tau\right\rangle,\left\langle 6,2,\tau\right\rangle,\left\langle 5,1,2,\tau\right\rangle,$ $\left\langle 5,1,4,\tau\right\rangle,$ $\left\langle 6,2,3,\tau\right\rangle\}$. Therefore, CF-CHAIN contains $7$ variables, one for each chain of length at most $3$.
HCF-CYCLE:
Suppose that $K=4$. If we do not change the ordering of the vertices, we have $\mathcal{H}=\{\left\langle 1,4\right\rangle,\left\langle 4,1\right\rangle,\left\langle 4,2\right\rangle,\left\langle 1,2,3\right\rangle,\left\langle 2,3,4\right\rangle,\left\langle 3,4,1\right\rangle\}$ and $\mathcal{H}_{1}=\emptyset$. Therefore, HCF-CYCLE contains $6$ variables, one for each half-cycle. For instance, cycle $\left\langle 1,2,3,4,1\right\rangle$ is obtained by setting $x_{\left\langle 1,2,3\right\rangle}=x_{\left\langle 3,4,1\right\rangle}=1$. We refer to Delorme et al., (2023) for an example that illustrates that HCF-CYCLE typically has fewer variables than CF-CYCLE for $K\geq 4$.
HCF-CHAIN:
Suppose that $L=4$. We have $\mathcal{H}^{\prime}_{\mathcal{N}}=\{\left\langle 5,1\right\rangle,\left\langle 5,4\right\rangle,\left\langle 6,2\right\rangle,\left\langle 5,1,2\right\rangle,\left\langle 5,1,4\right\rangle,$ $\left\langle 6,2,3\right\rangle\}$, $\mathcal{H}^{\prime}_{\tau}=\{\left\langle 1,\tau\right\rangle,\left\langle 2,\tau\right\rangle,\left\langle 4,\tau\right\rangle,\left\langle 1,2,\tau\right\rangle,\left\langle 1,4,\tau\right\rangle,\left\langle 2,3,\tau\right\rangle,\left\langle 3,4,\tau\right\rangle,\left\langle 4,1,\tau\right\rangle,\left\langle 4,2,\tau\right\rangle\}$ and $\mathcal{H}^{\prime}_{\mathcal{N}\tau}=\{\left\langle 5,\tau\right\rangle,$ $\left\langle 6,\tau\right\rangle\}$. Therefore, HCF-CHAIN contains $17$ variables, one for each half-chain and chain of length $1$. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{\left\langle 5,1,2\right\rangle}=y_{\left\langle 2,3,\tau\right\rangle}=1$.
EF-CYCLE:
Suppose that $K=3$. EF-CYCLE contains $6$ variables, one for each arc in $\mathcal{A}_{\mathcal{R}}$. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}=x_{34}=x_{42}=1$. Constraint (19) for maximal cycle-feasible path $\left\langle 1,2,3\right\rangle$ is required to forbid cycle $\left\langle 1,2,3,4,1\right\rangle$.
EF-CHAIN-EXP and EF-CHAIN-EXP:
Suppose that $L=4$. Arc set $\mathcal{A}$ contains $14$ arcs, resulting in $14$ variables for EF-CHAIN-EXP. EF-CHAIN-MTZ has $4$ additional timestamp variables, namely one per RDP. For instance, chains $\left\langle 6,2,3,4,\tau\right\rangle$ and $\left\langle 5,1,\tau\right\rangle$ are obtained by setting $y_{62}=y_{23}=y_{34}=y_{4\tau}=y_{51}=y_{1\tau}=1$, and for EF-CHAIN-MTZ also $t_{1}=1$, $t_{2}=1$, $t_{3}=2$ and $t_{4}=3$. Several long chains and cycles must be forbidden. For example, EF-CHAIN-EXP directly excludes chain $\left\langle 5,1,2,3,4,\tau\right\rangle$ through constraint (25) for minimal chain-infeasible path $\left\langle 1,2,3,4\right\rangle$, and it directly excludes cycle $\left\langle 2,3,4,2\right\rangle$ through constraint (26) corresponding to this cycle. On the other hand, EEF-CHAIN-EXP excludes these long chains and cycles indirectly via the timestamp variables.
EF-HYBRID:
Suppose that $K=2$ and $L=3$. Like EF-CHAIN-EXP, EF-HYBRID contains $14$ variables, one for each arc in $\mathcal{A}$. For instance, the solution consisting of cycle $\left\langle 1,4,1\right\rangle$ and chain $\left\langle 6,2,3,\tau\right\rangle$ is obtained by setting $z_{14}=z_{41}=z_{62}=z_{23}=z_{3\tau}=1$. Several long cycles and long chains must be forbidden. For example, cycle $\left\langle 2,3,4,2\right\rangle$ is excluded by the constraint (64) corresponding to this cycle, and chain $\left\langle 5,1,2,3,\tau\right\rangle$ is excluded by constraint (65) for minimal chain-infeasible path $\left\langle 1,2,3\right\rangle$.
EEF-CYCLE:
Suppose that $K=4$. The subgraphs $\mathcal{G}^{1}$, $\mathcal{G}^{2}$, $\mathcal{G}^{3}$ and $\mathcal{G}^{4}$ required by EEF-CYCLE are presented in Figure 4. Overall, there are $10$ arcs across the subgraphs, leading to $10$ variables. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}^{2}=x_{34}^{2}=x_{42}^{2}=1$. Note that even though these arcs are present in subgraph $\mathcal{G}^{1}$ too, any solution with $x_{23}^{1}=x_{34}^{1}=x_{42}^{1}=1$ is forbidden due to constraints (34), as no arc leaving vertex $1$ is selected in that subgraph.
1234
1234
1234
1234
EEF-CHAIN-EXP and EEF-CHAIN-MTZ:
Suppose that $L=4$. The subgraphs $\mathcal{G}^{5}$ and $\mathcal{G}^{6}$ required by EEF-CHAIN-EXP and EEF-CHAIN-MTZ are presented in Figure 5. The total number of arcs across both subgraphs is $24$, leading to $24$ variables for EEF-CHAIN-EXP. EEF-CHAIN-MTZ has $4$ additional timestamp variables, namely one per RDP. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{51}^{5}=y_{12}^{5}=y_{23}^{5}=y_{3\tau}=1$ and for EEF-CHAIN-MTZ also $t_{1}=1$, $t_{2}=2$, $t_{3}=3$ and $t_{4}\in[1,3]$. Note that chain $\left\langle 6,2,3,4,1,\tau\right\rangle$ is infeasible due to constraints (49), as this chain would require too many arcs to be selected in subgraph $\mathcal{G}^{6}$. Moreover, cycle $\left\langle 1,4,1\right\rangle$ is excluded directly for EEF-CHAIN-EXP due to the constraint (50) corresponding to this cycle, while it is excluded indirectly for EEF-CHAIN-MTZ via the timestamp variables.
[…]
PIEF-CYCLE:
Suppose that $K=4$. The subgraphs required by PIEF-CYCLE are depicted in Figure 6, where the labels on the arcs are the sets $\mathcal{K}^{s}(u,v)$ of possible positions of the arcs in the subgraphs. In total there are $9$ variables, one for each feasible position of each arc in each subgraph. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}^{21}=x_{34}^{22}=x_{42}^{23}=1$.
[…]
1234
PIEF-CHAIN:
Suppose that $L=4$. The graph required by PIEF-CHAIN is depicted in Figure 7, where the labels on the arcs are the sets $\mathcal{K}^{\prime}(u,v)$ of possible positions of the arcs. In total there are $20$ variables, one for each feasible position of each arc. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{51}^{1}=y_{12}^{2}=y_{23}^{3}=y_{3\tau}^{4}=1$.
[…]
Appendix E Summary of model properties
@@
Cycles CF-CYCLE $O(|\mathcal{R}|^{K})$ $O(|\mathcal{R}|)$ -
HCF-CYCLE $O(|\mathcal{R}|^{1+\left\lceil K/2\right\rceil})$ $O(|\mathcal{R}|^{2})$ -
@@
EEF-CYCLE $O(|\mathcal{R}||\mathcal{A}_{\mathcal{R}}|)$ $O(|\mathcal{R}|^{2})$ -
PIEF-CYCLE $O(K|\mathcal{R}||\mathcal{A}_{\mathcal{R}}|)$ $O(K|\mathcal{R}|^{2})$ -
Chains CF-CHAIN $O(|\mathcal{N}||\mathcal{R}|^{L-1})$ $O(|\mathcal{N}|+|\mathcal{R}|)$ -
HCF-CHAIN $O(|\mathcal{N}||\mathcal{R}|^{\left\lfloor L/2\right\rfloor}+|\mathcal{R}|^{\left\lceil L/2\right\rceil})$ $O(|\mathcal{N}|+|\mathcal{R}|)$ -
@@
EF-CHAIN-MTZ $O(|\mathcal{A}|)$ $O(|\mathcal{N}|+|\mathcal{A}_{\mathcal{R}}|)$ -
@@
EEF-CHAIN-MTZ $O(|\mathcal{N}||\mathcal{A}|)$ $O(|\mathcal{N}||\mathcal{R}|+|\mathcal{A}_{\mathcal{R}}|)$ -
PIEF-CHAIN $O(L|\mathcal{A}|)$ $O(|\mathcal{N}|+L|\mathcal{R}|)$ -
@@
Appendix F Graph Reduction Algorithms
@@
EF-CYCLE $\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:\exists u\in\mathcal{R}\text{ s.t. }(u,v)\in\mathcal{A}_{\mathcal{R}},d_{vu}+1\leq K\}$
@@
EF-CHAIN-EXP/MTZ $\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:d_{\mathcal{N}v}+1\leq L\}$
@@
EF-HYBRID $\widetilde{\mathcal{R}}=\{v\in\mathcal{R}:(\exists u\in\mathcal{R}\text{ s.t. }(u,v)\in\mathcal{A}_{\mathcal{R}},d_{vu}+1\leq K)\vee(d_{\mathcal{N}v}+1\leq L)\}$
@@
[…]
@@
Instead, we propose to construct the sets $\mathcal{K}^{s}(u,v)$ by running Algorithm 1 once for each subgraph $\mathcal{G}^{s}$ for $s\in\mathcal{R}$. Note that every run is essentially a breadth-first search of the considered subgraph $\mathcal{G}^{s}$, starting from vertex $s$, where a position $k$ is added to the set of possible positions for an arc if (i) the tail of the arc is the head of at least one arc for which position k1k-1 is possible, and (ii) the cycle can still be closed using at most KkK-k arcs.
@@
[…]
@@
[…]
@@
Appendix G Implicitly Dealing with the Terminal Vertex